# Cyber Attacks Detection, Prevention, and Source Localization in Digital Substation Communication using Hybrid Statistical-Deep Learning

### v1

@@ VI Conclusions
This paper presents a novel hybrid statistical-deep learning-based method for the detection, prevention, and attack source location of IEC 61850 SV injection attacks in digital substations. It has been shown how the usage of SV frames arrival time statistics provides useful information to discern between legitimate and malicious SV frames, early detected advanced cyber attacks, e.g., MitM, and identify malicious IEDs within the digital substation communication network. As validated by the experimental results, the proposed method provides promising intrusion detection and prevention performances while guaranteeing no interference to protection schemes when no cyber attack is ongoing. Furthermore, it meets IEC 61850-5 and IEC 61869-9 strict latency requirements and required throughput, respectively. The method is robust to communication network latency variations and time synchronization issues. Thus, it provides a viable solution to enhance cyber security and resilience in digital substations. Although the method was validated on IEC 61850-compliant testbeds, future work consists in the further engineering of the proposed method and its testing in real digital substations. Moreover, advanced attacker techniques capable of maximizing the likelihood of malicious SV frames acceptance, and how to defend against such advanced threats, will be investigated.

### v2

@@ V Conclusions
This paper presents a novel integrated hybrid statistical-deep learning-based method for detecting, preventing, and localizing IEC 61850 SV-based cyber attacks in digital substations. It has been shown that the statistical modeling of SV frames arrival time combined with DL enables the distinction between legitimate and malicious SV frames, early detection of MitM attacks on SV streams, and localization of malicious IEDs within the digital substation communication network. Experimental validation demonstrates promising intrusion detection, prevention, and localization performances while ensuring no interference to protection schemes in the absence of cyber attacks. The method satisfies IEC 61850-5 and IEC 61869-9 strict latency requirements and required throughput, respectively, and is robust to communication network latency, jitter, and time synchronization deviations. Thus, it provides a viable solution to enhance cyber security and resilience in digital substations. Although validated on IEC 61850-compliant testbeds, future work will focus on field deployment in operational digital substations and defenses against advanced attacker strategies designed to evade frame acceptance filtering.